# Supersymmetry and Localization on Three-Dimensional Orbifolds

### v1

@@ 2.4 The squashed lens space
2.5 The branched, squashed three-sphere
Supersymmetric gauge theories on branched spheres, namely branched coverings of $S^{3}$, were explored in [32] because of their connection to Rényi entropy as well as entanglement entropy. Analogous results for free theories on branched spheres were previously obtained in [45]. Here we consider supersymmetric gauge theories on a more general orbifold, namely a branched squashed three-sphere $S^{3}_{q,p,{n_{+}},{n_{-}}}$, whose line element reads
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+\frac{q^{2}}{{n_{+}}^{2}}\sin^{2}x\,{\rm d}\varphi^{2}+\frac{p^{2}}{{n_{-}}^{2}}\cos^{2}x\,{\rm d}\psi^{2}~{},$ (2.78)
where $\varphi,\psi\in{\left[{0,2\pi}\right]}$ parametrize the two circles $S^{1}$ in $S^{3}_{q,p,{n_{+}},{n_{-}}}$, while $x\in{\left[{0,\pi/2}\right]}$. The (a priori complex) numbers $q,p$ are the squashing parameters also appearing in the function
$f=-\sqrt{p^{2}\sin^{2}x+q^{2}\cos^{2}x}~{}.$ (2.79)
The positive integers ${\left({{n_{+}},{n_{-}}}\right)}$ parametrize the orbifold singularities at $x=0$ and $x=\pi/2$. If ${n_{+}}={n_{-}}=1$, we obtain the squashed sphere studied in [12]. Instead, the metric of the branched sphere considered in [32] is obtained from (2.78) after choosing $q=p=1$ and identifying ${n_{+}}=q^{-1}_{\rm there}$ as well as ${n_{-}}=p^{-1}_{\rm there}$. Further setting ${n_{+}}=1$ yields the line element of the orbifold employed in [45].
@@ 2.5 The branched, squashed three-sphere
$e^{2}$ $=q\sin x\,{\rm d}\varphi/{n_{+}}~{},$
$e^{3}$ $=p\cos x\,{\rm d}\psi/{n_{-}}~{}.$ (2.80)
The branched,squashed three-sphere is linked to the other orbifold backgrounds explored in this paper as $S^{3}_{q,p,{n_{+}},{n_{-}}}$ is a non-trivial $S^{1}$ fibration with base a spindle $\mathbb{\Sigma}$. Indeed, by applying a $SL{\left({2,\mathbb{Z}}\right)}$ map acting on ${\left({\varphi,\psi}\right)}$ as
$\varphi=t_{11}\widehat{\varphi}_{1}+t_{12}\widehat{\varphi}_{2}~{},\qquad\psi=t_{21}\widehat{\varphi}_{1}+t_{22}\widehat{\varphi}_{2}~{},$ $t_{11}t_{22}-t_{12}t_{21}=1~{},$ (2.81)
@@
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+c_{22}{\rm d}\widehat{\varphi}_{1}^{2}+h{\left({{\rm d}\widehat{\varphi}_{2}+{A^{(1)}}}\right)}^{2}~{},$ (2.82)
where $\widehat{\varphi}_{2}\in{\left[{0,2\pi}\right]}$ parametrizes $S^{1}$ while $x\in{\left[{0,\pi/2}\right]}$ and $u_{1}\in{\left[{0,2\pi}\right]}$ are coordinates on $\mathbb{\Sigma}$. The objects appearing in the line element immediately above read
@@
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}x{\rm d}\widehat{\varphi}_{1}\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=\frac{1}{{n_{-}}t_{12}}+\frac{1}{{n_{+}}t_{22}}=\chi_{\rm orb}~{},$ (2.84)
confirming that the branched squashed three-sphere is a $S^{1}$ fibration over the base encoded by the spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{{n_{+}}t_{22},{n_{-}}t_{12}}\right]}}$. Moreover,the flux of $A_{1}$ through $\mathbb{\Sigma}$ is non-vanishing and fractional:
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{A^{(1)}}=\frac{t_{11}}{t_{12}}-\frac{t_{21}}{t_{22}}=\frac{{n_{+}}{n_{-}}}{{\left({{n_{+}}t_{22}}\right)}{\left({{n_{-}}t_{12}}\right)}}~{},$ (2.85)
verifying that the fibration is non-trivial and that the background field $A_{1}$ is a 1-form representing a $\mathcal{O}(-{n_{+}}{n_{-}})$ orbibundle on $\mathbb{\Sigma}$.
@@
$\zeta_{\alpha}$ $=\sqrt{\frac{\mathcal{N}p}{2\,{n_{-}}}}\,e^{\frac{{\rm i}}{2}{\left({\alpha_{2}\varphi+\alpha_{3}\psi}\right)}}\sqrt{1-\mathfrak{s}_{w}\sin x}\begin{pmatrix}1\\ \sec x+\mathfrak{s}_{w}\tan x\end{pmatrix}_{\alpha}~{},$ (2.88)
$\widetilde{\zeta}_{\alpha}$ $=-\sqrt{\frac{\mathcal{N}p}{2\,{n_{-}}}}\,e^{-\frac{{\rm i}}{2}{\left({\alpha_{2}\varphi+\alpha_{3}\psi}\right)}}\sqrt{1-\mathfrak{s}_{w}\sin x}\begin{pmatrix}\sec x+\mathfrak{s}_{w}\tan x\\ 1\end{pmatrix}_{\alpha}~{},$ (2.91)
@@
$A^{C}$ $=\frac{1}{2}{\left({-\frac{q}{{n_{+}}f}+\alpha_{2}}\right)}{\rm d}\varphi+\frac{1}{2}{\left({-\frac{\mathfrak{s}_{w}p}{{n_{-}}f}+\alpha_{3}}\right)}{\rm d}\psi~{},$ (2.92)
is turned on. The integrability conditions for this background are verified for any $\alpha_{2,3}$, even though the 1-form $A^{C}$ vanishes at the conical points of the branched, squashed three-sphere only if
$\alpha_{2}=-1/{n_{+}}~{},$ $\alpha_{3}=-\mathfrak{s}_{w}/{n_{-}}~{}.$ (2.93)
@@
$V$ $={\left({\frac{1}{f}+{\rm i}\,\mathfrak{s}_{w}H}\right)}{\left({\frac{q}{{n_{+}}}\sin^{2}x\,{\rm d}\varphi+\mathfrak{s}_{w}\frac{\,p}{{n_{-}}}\cos^{2}x\,{\rm d}\psi}\right)}~{},$

### v2

@@ 2.4 The squashed lens space
2.5 The branched-squashed three-sphere
Supersymmetric gauge theories on branched spheres, namely branched coverings of $S^{3}$, were explored in [33] because of their connection to Rényi entropy as well as entanglement entropy. Analogous results for free theories on branched spheres were previously obtained in [46]. Here we consider supersymmetric gauge theories on a more general orbifold, namely a branched squashed three-sphere $S^{3}_{b_{1},b_{2},p_{+},p_{-}}$, whose line element reads
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+\frac{b_{1}^{2}}{p_{+}^{2}}\sin^{2}x\,{\rm d}\varphi_{1}^{2}+\frac{b_{2}^{2}}{p_{-}^{2}}\cos^{2}x\,{\rm d}\varphi_{2}^{2}~{},$ (2.68)
where $\varphi_{1},\varphi_{2}\in{\left[{0,2\pi}\right]}$ parametrize the two circles in $S^{3}_{b_{1},b_{2},p_{+},p_{-}}$, while $x\in{\left[{0,\pi/2}\right]}$. The (a priori complex) numbers $b_{1},b_{2}$ are the squashing parameters also appearing in the function
$f=-\sqrt{b_{1}^{2}\cos^{2}x+b_{2}^{2}\sin^{2}x}~{}.$ (2.69)
The positive integers ${\left({p_{+},p_{-}}\right)}$ parametrize the orbifold singularities at $x=0$ and $x=\pi/2$. If $p_{+}=p_{-}=1$, we obtain the squashed sphere studied in [11]. Instead, the metric of the branched sphere considered in [33] is obtained from (2.68) after choosing $b_{1}=b_{2}=1$ and identifying $p_{+}=q^{-1}_{\rm there}$ as well as $p_{-}=p^{-1}_{\rm there}$. Further setting $p_{+}=1$ yields the line element of the orbifold employed in [46].
@@ 2.5 The branched-squashed three-sphere
$e^{2}$ $=b_{1}\sin x\,{\rm d}\varphi_{1}/p_{+}~{},$
$e^{3}$ $=b_{2}\cos x\,{\rm d}\varphi_{2}/p_{-}~{}.$ (2.70)
The branched,squashed three-sphere is linked to the other orbifold backgrounds explored in this paper as $S^{3}_{b_{1},b_{2},p_{+},p_{-}}$ is a non-trivial $S^{1}$ fibration with base a spindle $\mathbb{\Sigma}$. Indeed, by applying a $SL{\left({2,\mathbb{Z}}\right)}$ map acting on ${\left({\varphi_{1},\varphi_{2}}\right)}$ as
$\varphi_{1}=t_{11}\varphi+t_{12}\psi~{},\qquad\varphi_{2}=t_{21}\varphi+t_{22}\psi~{},$ $t_{11}t_{22}-t_{12}t_{21}=1~{},$ (2.71)
@@
${\rm d}s^{2}=f^{2}{\rm d}x^{2}+c_{22}{\rm d}\varphi^{2}+h{\left({{\rm d}\psi+{A^{(1)}}}\right)}^{2}~{},$ (2.72)
where $\psi\in{\left[{0,2\pi}\right]}$ parametrizes the $S^{1}$ fiber while $x\in{\left[{0,\pi/2}\right]}$ and $\varphi\in{\left[{0,2\pi}\right]}$ are coordinates on the base $\mathbb{\Sigma}$. The objects appearing in the line element (2.72) read
@@
$\frac{1}{4\pi}\int_{\mathbb{\Sigma}}{\rm d}x{\rm d}\varphi\sqrt{g_{\mathbb{\Sigma}}}R_{\mathbb{\Sigma}}=\frac{1}{p_{-}t_{12}}+\frac{1}{p_{+}t_{22}}=\chi_{\rm orb}~{},$ (2.74)
confirming that the branched squashed three-sphere is a $S^{1}$ fibration over the base encoded by the spindle $\mathbb{\Sigma}=\mathbb{WCP}^{1}_{{\left[{p_{+}t_{22},p_{-}t_{12}}\right]}}$. Especially, the flux of ${A^{(1)}}$ through $\mathbb{\Sigma}$ is non-vanishing and fractional:
$\frac{1}{2\pi}\int_{\mathbb{\Sigma}}{\rm d}{A^{(1)}}=\frac{t_{11}}{t_{12}}-\frac{t_{21}}{t_{22}}=\frac{p_{+}t_{11}}{p_{+}t_{12}}-\frac{p_{-}t_{21}}{p_{-}t_{22}}=\frac{p_{+}p_{-}}{{\left({p_{+}t_{22}}\right)}{\left({p_{-}t_{12}}\right)}}~{},$ (2.75)
verifying that the fibration is non-trivial and that the background field $A^{(1)}$ is a 1-form representing the orbibundle $\mathcal{O}(-p_{+}p_{-})$ on $\mathbb{\Sigma}$. The line element (2.72) in a neighborhood of $x=0$ reads
${\rm d}s^{2}|_{x=0}=b_{1}^{2}{\left[{{\rm d}x^{2}+x^{2}{\left({\frac{{\rm d}\varphi}{p_{+}t_{22}}}\right)}^{2}}\right]}+{\left({\frac{b_{2}t_{22}}{p_{-}}}\right)}^{2}{\left({{\rm d}\psi+\frac{t_{21}}{t_{22}}{\rm d}\varphi}\right)}^{2}~{},$ (2.76)
which can be covered by the smooth manifold $\mathbb{C}\times\mathbb{C}^{*}$ by defining the complex coordinates
$z_{0}^{(1)}=b_{1}x\exp{\left({\frac{{\rm i}\varphi}{p_{+}t_{22}}}\right)}~{},$ $z_{0}^{(2)}=\frac{b_{2}t_{22}}{p_{-}}\exp{\left[{{\rm i}{\left({\psi+\frac{t_{21}}{t_{22}}\varphi}\right)}}\right]}~{},$ (2.77)
where $z_{0}^{(2)}\in\mathbb{C}^{*}$ embeds $S^{1}$ in $\mathbb{C}$ as $|z_{0}^{(2)}|^{2}=|{b_{2}t_{22}}/{p_{-}}|^{2}$ is constant. Moreover, $z_{0}^{(2)}$ is invariant under $\psi\sim\psi+2\pi$, while the rotational invariance of the base provides the identifications
$\varphi\sim\varphi+2\pi\implies{\left({z_{0}^{(1)},z_{0}^{(2)}}\right)}\sim{\left({w_{+}z_{0}^{(1)},w_{+}^{p_{+}t_{21}}z_{0}^{(2)}}\right)}~{},$ $w_{+}=\exp{\left({\frac{2\pi{\rm i}}{p_{+}t_{22}}}\right)}~{}.$ (2.78)
In particular, going $t_{22}$ times around the point $x=0$ gives
$\varphi\sim\varphi+2\pi\,t_{22}\implies{\left({z_{0}^{(1)},z_{0}^{(2)}}\right)}\sim{\left({w_{+}^{t_{22}}z_{0}^{(1)},z_{0}^{(2)}}\right)}={\left({e^{2\pi{\rm i}/p_{+}}z_{0}^{(1)},z_{0}^{(2)}}\right)}~{},$ (2.79)
indicating that $x=0$ is the fixed point of a non-free $\mathbb{Z}_{p_{+}}$ action affecting the base of the fibration but preserving the fiber. This confirmes that the branched squashed sphere displays a genuine conical singularity at $x=0$ encoded by the quotient $\mathbb{C}/\mathbb{Z}_{p_{+}}$. Analogously, near $x=\pi/2$ we find
${\rm d}s^{2}|_{x=\pi/2}=b_{2}^{2}{\left[{{\rm d}x^{2}+{\left({\frac{\pi}{2}-x}\right)}^{2}{\left({\frac{{\rm d}\varphi}{p_{-}t_{12}}}\right)}^{2}}\right]}+{\left({\frac{b_{1}t_{12}}{p_{+}}}\right)}^{2}{\left({{\rm d}\psi+\frac{t_{11}}{t_{12}}{\rm d}\varphi}\right)}^{2}~{},$ (2.80)
which can be lifted to the smooth manifold $\mathbb{C}\times\mathbb{C}^{*}$ after introducing the complex coordinates
$z_{\pi/2}^{(1)}=b_{2}{\left({\frac{\pi}{2}-x}\right)}\exp{\left({\frac{-{\rm i}\varphi}{p_{-}t_{12}}}\right)}~{},$ $z_{\pi/2}^{(2)}=\frac{b_{1}t_{12}}{p_{+}}\exp{\left[{{\rm i}{\left({\psi+\frac{t_{11}}{t_{12}}\varphi}\right)}}\right]}~{},$ (2.81)
with $z_{\pi/2}^{(2)}\in\mathbb{C}^{*}$ embedding $S^{1}$ into $\mathbb{C}$. Alike $z_{0}^{(2)}$, the complex coordinate $z_{\pi/2}^{(2)}$ is not affected by shifts $\psi\to\psi+2\pi$, whereas
$\varphi\sim\varphi+2\pi\implies{\left({z_{\pi/2}^{(1)},z_{\pi/2}^{(2)}}\right)}\sim{\left({w_{-}^{-1}z_{\pi/2}^{(1)},w_{-}^{-p_{-}t_{11}}z_{\pi/2}^{(2)}}\right)}~{},$ $w_{-}=\exp{\left({\frac{2\pi{\rm i}}{p_{-}t_{12}}}\right)}~{}.$ (2.82)
Winding $t_{12}$ times around the point $x=\pi/2$ yields
$\varphi\sim\varphi+2\pi\,t_{12}\implies{\left({z_{\pi/2}^{(1)},z_{\pi/2}^{(2)}}\right)}\sim{\left({w_{-}^{-t_{12}}z_{\pi/2}^{(1)},z_{\pi/2}^{(2)}}\right)}={\left({e^{-2\pi{\rm i}/p_{-}}z_{\pi/2}^{(1)},z_{\pi/2}^{(2)}}\right)}~{},$ (2.83)
making manifest that $\mathbb{Z}_{p_{-}}$ does not act freely on the base at $x=\pi/2$ and leaves the fiber invariant. Thus, the branched squashed sphere exhibits a $\mathbb{C}/\mathbb{Z}_{p_{-}}$ conical singularity at $x=\pi/2$.
@@
$\zeta_{\alpha}$ $=\sqrt{\frac{k_{0}b_{2}}{2\,p_{-}}}\,e^{\frac{{\rm i}}{2}{\left({\alpha_{2}\varphi_{1}+\alpha_{3}\varphi_{2}}\right)}}\sqrt{1-\mathfrak{s}_{w}\sin x}\begin{pmatrix}1\\ \sec x+\mathfrak{s}_{w}\tan x\end{pmatrix}_{\alpha}~{},$
$\widetilde{\zeta}_{\alpha}$ $=-\sqrt{\frac{k_{0}b_{2}}{2\,p_{-}}}\,e^{-\frac{{\rm i}}{2}{\left({\alpha_{2}\varphi_{1}+\alpha_{3}\varphi_{2}}\right)}}\sqrt{1-\mathfrak{s}_{w}\sin x}\begin{pmatrix}\sec x+\mathfrak{s}_{w}\tan x\\ 1\end{pmatrix}_{\alpha}~{},$ (2.84)
@@
$A^{C}$ $=\frac{1}{2}{\left({-\frac{b_{1}}{p_{+}f}+\alpha_{2}}\right)}{\rm d}\varphi_{1}+\frac{1}{2}{\left({-\frac{\mathfrak{s}_{w}b_{2}}{p_{-}f}+\alpha_{3}}\right)}{\rm d}\varphi_{2}~{},$ (2.85)
is turned on. The integrability conditions for this background are verified for any $\alpha_{2,3}$, even though the 1-form $A^{C}$ vanishes at the conical points of the branched-squashed three-sphere only if
$\alpha_{2}=-1/p_{+}~{},$ $\alpha_{3}=-\mathfrak{s}_{w}/p_{-}~{}.$ (2.86)
@@
$V$ $={\left({\frac{1}{f}+{\rm i}\,\mathfrak{s}_{w}H}\right)}{\left({\frac{b_{1}}{p_{+}}\sin^{2}x\,{\rm d}\varphi_{1}+\mathfrak{s}_{w}\frac{b_{2}}{p_{-}}\cos^{2}x\,{\rm d}\varphi_{2}}\right)}~{},$